how long does it take for a deposit of $900 to double at 10% compounded continuously
Answer by stanbon(75887)

how long does it take for a deposit of $900 to double at 10% compounded continuously
-----------
A(t) = P*e^(rt)
----
1800 = 900*e^(0.10t)
----
e^(0.1t) = 2
---
0.1t = ln2
----
t = 6.93 years
====================
Cheers,
Stan H.
====================